# The computational power of discrete chemical reaction networks with bounded executions

### v2

@@ 6 Limitations of execution bounded CRNs
The main positive results of the paper (LABEL:{thm:execution_bdd_CRDs_decide_all_semilinear_sets} and 5.5) rely on the assumption that valid initial configurations have a single leader (in particular, they are execution bounded only from configurations with a single leader, but not from arbitrary configurations). Section 4 shows that we may assume the CRD deciding a semilinear set is all-voting. However, for the “constructive” results LABEL:{lem:boolean-closure-crd} and 5.5, which compose the output of a CRD $\mathcal{D}$ with downstream computation, using $\mathcal{D}$ as a “subroutine” to stably compute a more complex set/function, the constructions crucially use the assumption that $\mathcal{D}$ is single-voting (i.e., only the leader of $\mathcal{D}$ votes) to argue the resulting composed CRN is execution bounded. In this section we show these assumptions are necessary, proving that execution bounded CRNs without those constraints are severely limited in their computational abilities.
We show that entirely execution bounded CRNs (from every configuration) can be characterized by a simpler property of having a “linear potential function” that essentially measures how close the CRN is to reaching a terminal configuration. We then use this characterization to prove that entirely execution bounded CRNs can stably decide only limited semilinear predicates (eventually constant, Definition 6.11), assuming all species vote, and that molecular counts cannot decrease to $O(1)$ in stable configurations (see Definition 6.8).
@@ 6.1 Linear potential functions
We define a linear potential function of a CRN to be a nonnegative linear function of states that each reaction strictly decreases.
@@
A linear potential function $\Phi:\mathbb{R}^{\Lambda}\to\mathbb{R}_{\geq 0}$ for a CRN is a nonnegative, linear function of configurations, such that for each reaction $(\mathbf{r},\mathbf{p})$, $\Phi(\mathbf{p}-\mathbf{r})<0.$
Note that since $\Phi(\mathbf{x})=\sum_{S\in\Lambda}c_{S}\mathbf{x}(S)$ is required to be nonnegative on all configurations $\mathbf{x}$, it must be nondecreasing in each species, i.e., all coefficients $c_{S}$ must be nonnegative (though some are permitted to be 0). Intuitively, we can think of $\Phi$ as assigning a nonnegative “mass” to each species (the mass of $S$ is $c_{S}$), such that each reaction removes a positive amount of mass from the system.
@@
A system of linear inequalities with rational coefficients has a real solution if and only if it has a rational solution. For any homogeneous system (where all inequalities are comparing to 0), any positive scalar multiple of a solution is also a solution. By clearing denominators, a system has a rational solution if and only if it has an integer solution. Thus, one can equivalently define a linear potential function to be a function $\Phi(\mathbf{x})=\sum_{S\in\Lambda}c_{S}\mathbf{x}(S)$ such that each $c_{S}\in\mathbb{N}$, i.e., we may assume $\Phi:\mathbb{N}^{\Lambda}\to\mathbb{N}$. In particular, since $\Phi$ is decreased by each reaction, it is decreased by at least 1.
@@
$(\forall i)\sum_{S\in\Lambda}[\mathbf{p}_{i}(S)-\mathbf{r}_{i}(S)]c_{S}<0$
[…]
submission,final $(\forall i)\sum_{S\in\Lambda}[\mathbf{p}_{i}(S)-\mathbf{r}_{i}(S)]c_{S}<0.$ For example, for the reactions $A+A\mathop{\rightarrow}\limits B+C$ and $B+B\mathop{\rightarrow}\limits A$, for each reaction to strictly decrease the potential function $\Phi(\mathbf{x})=c_{A}\mathbf{x}(A)+c_{B}\mathbf{x}(B)+c_{C}\mathbf{x}(C)$, $\Phi$ must satisfy $2c_{A}>c_{B}+c_{C}$ and $2c_{B}>c_{A}$. In this case, $c_{A}=1,c_{B}=1,c_{C}=0$ works.
The following is a variant of Farkas’ Lemma [farkas1902theorie], one of several similar “Theorems of the Alternative” stating that exactly one of two different linear systems has a solution. (See [mangasarian1994nonlinear, Chapter 4] for a list of such theorems.) A proof can be found in [gale1960theory, Theorem 2.10].
@@
Although we do not need the following fact, it is worthwhile to observe that, if $\mathbf{M}$ is integer-valued (as in our application), then the solution $\mathbf{u}$ or $\mathbf{v}$ (whichever exists) in Section 6.1 has entries that are at most exponential in $\|\mathbf{M}\|$ i.e., at most exponential in the sum of absolute values of entries of $\mathbf{M}$ (see e.g., [papadimitriou1981complexity]). So in particular when we consider $\mathbf{M}$ having small $O(1)$ size entries, this means the solution $\mathbf{u}$ or $\mathbf{v}$ has entries that are at most exponential in the number of rows and columns of $\mathbf{M}$. When $\mathbf{M}$ is a stoichiometric matrix, this corresponds to the number of species and reactions, respectively, of the CRN.
@@
Theorem 6.5.
@@
Let $\mathcal{C}=(\Lambda,R)$ be a CRN. The forward direction is easy: assuming $\mathcal{C}$ has potential function $\Phi$, since each reaction decreases $\Phi$ by at least 1 (see Remark 6.2), starting from configuration $\mathbf{x}$, we can execute at most $\Phi(\mathbf{x})$ reactions while keeping $\Phi$ nonnegative. Thus $\mathcal{C}$ is entirely execution bounded.
@@
By Section 6.1, there is an integer vector $\mathbf{v}\geqq\mathbf{0}$ such that $\mathbf{v}\mathbf{M}<\mathbf{0}$. Let $\mathbf{v}\in\mathbb{N}^{\Lambda}$ be the coefficients of a linear function $\Phi:\mathbb{N}^{\Lambda}\to\mathbb{N}$, i.e., $\Phi(\mathbf{x})=\mathbf{v}\cdot\mathbf{x}$. Then the vector $\mathbf{v}\mathbf{M}\in\mathbb{Z}^{R}$ represents the amount $\Phi$ changes by one unit of each reaction, i.e., $\mathbf{v}\mathbf{M}(\alpha)$ is the amount $\Phi$ increases when executing reaction $\alpha$ once. Since $\mathbf{v}\mathbf{M}<\mathbf{0}$, this means that every reaction strictly decreases $\Phi$, i.e., $\Phi$ is a linear potential function for $\mathcal{D}$.
@@
By employing the real-valued version of Section 6.1, the above proof also shows that Theorem 6.5 holds for the continuous model of CRNs [chen2023rate], in which species amounts are modeled as continuous nonnegative real concentrations. In this case, a continuous CRN would be defined to be execution bounded from configuration $\mathbf{x}$ if each reaction can be executed by at most a finite (real-valued) amount from $\mathbf{x}$.
@@ 6.2 Impossibility of stably deciding majority and parity
No noncollapsing, all-voting, entirely execution bounded CRD can stably decide the majority predicate $[X_{1}\geq X_{2}?]$ or the parity predicate $[X\equiv 1\mod 2?]$.
@@
Continuing in this way, we have a sequence of stable configurations $\mathbf{a}_{1},\mathbf{b}_{1},\mathbf{a}_{2},\mathbf{b}_{2},\dots$ where all species in $\mathbf{a}_{i}$ vote yes and all species in $\mathbf{b}_{i}$ vote no. Since $\mathcal{D}$ is noncollapsing, the size of the configurations $\mathbf{a}_{i}$ and $\mathbf{b}_{i}$ increases without bound as $i\to\infty$. (Possibly $\|\mathbf{a}_{i+1}\|<\|\mathbf{a}_{i}\|$, i.e., the size is not necessarily monotonically increasing, but for all sufficiently large $j>i$, we have $\|\mathbf{a}_{j}\|>\|\mathbf{a}_{i}\|$.) Since all species vote, for some constant $\delta>0$, to get from $\mathbf{a}_{i}+\{X_{2}\}$ to $\mathbf{b}_{i}$, at least $\delta\|\mathbf{a}_{i}\|$ reactions must occur. This is because all species in $\mathbf{a}_{i}$ must be removed since they vote yes, and each reaction removes at most $O(1)$ molecules. (Concretely, let $\delta=1/\max_{(\mathbf{r},\mathbf{p})\in R}\|\mathbf{r}\|-\|\mathbf{p}\|$, i.e., 1 over the most net molecules consumed in any reaction.) Similarly, to get from $\mathbf{b}_{i}+\{X_{1}\}$ to $\mathbf{a}_{i+1}$, at least $\delta\|\mathbf{b}_{i}\|$ reactions must occur.
Since $\mathcal{D}$ is entirely execution bounded, by Theorem 6.5, $\mathcal{D}$ has a linear potential function $\Phi(\mathbf{x})=\mathbf{v}\cdot\mathbf{x}$, where $\mathbf{v}\geq\mathbf{0}$. Adding a single $X_{2}$ to $\mathbf{a}_{i}$ increases $\Phi$ by the constant $\mathbf{v}(X_{2})$. Since $\|\mathbf{a}_{i}\|$ grows without bound, the number of reactions to get from $\mathbf{a}_{i}+\{X_{2}\}$ to $\mathbf{b}_{i}$ increases without bound as $i\to\infty$, and since each reaction strictly decreases $\Phi$ by at least 1, the total change in $\Phi$ that results from adding $X_{2}$ and then going from $\mathbf{a}_{i}+\{X_{2}\}$ to $\mathbf{b}_{i}$ is unbounded in $i$, so unboundedly negative for sufficiently large $i$ (negative once $i$ is large enough that $\delta\|\mathbf{a}_{i}\|\geq\mathbf{v}(X_{2})+2$). Similarly, adding a single $X_{1}$ to $\mathbf{b}_{i}$ and going from $\mathbf{b}_{i}+\{X_{1}\}$ to $\mathbf{a}_{i+1}$, the resulting total change in $\Phi$ is unbounded and (eventually) negative.
@@
However, since $\mathcal{D}$ is noncollapsing, this can only happen for a constant number of $i$ (so $\Phi$ never reaches more than a constant above its initial value $\Phi(\mathbf{s}+\{X_{1},X_{2}\})$), after which $\Phi$ strictly decreases after each round of this process. At some point in this process, $\mathcal{D}$ will not be able to reach all the way to the next $\mathbf{a}_{i}$ or $\mathbf{b}_{i}$ without $\Phi$ becoming negative, a contradiction.
@@
Theorem 6.9 is false without the noncollapsing hypothesis. The following collapsing, leaderless (but all-voting and entirely execution bounded) CRD stably decides majority: Species $X_{1},x_{1}$ vote yes, while $X_{2},x_{2}$ vote no:
@@
It has bounded executions from any configuration: $\min(\#X_{1},\#X_{2})$ of the first reaction can occur, and the other reactions decrease molecular count, so are limited by the total configuration size. However, it is collapsing since a stable configuration of size 1 is always reachable. Theorem 6.9 is similarly false without the all-voting hypothesis; for each of the reactions with one product above, add another non-voting product $W$. This converts the CRD to be noncollapsing but not all-voting. Of course, the execution bounded hypothesis is also necessary: the original population protocols paper [angluin2004computation] showed that all-voting, noncollapsing, leaderless population protocols can stably decide all semilinear predicates.
@@
full It has bounded executions from any configuration: exactly #$\#X_{1}+\#X_{0}-1$ reactions can occur since each reduces #$\#X_{1}+\#X_{0}$ by 1. Similar to above, by adding the non-voting product $W$ to each reaction above, this CRD becomes noncollapsing but not all-voting, showing that the all-voting hypothesis is also necessary for stably deciding parity.
@@ 6.3 Impossibility of stably deciding not eventually constant predicates
[…]
@@
In other words, $A$ is periodic if it is a union of copies of a finite subset $F$ of the $k\times k\times\dots\times k$ hypercube with a corner at the origin, translated in each direction by every nonnegative integer multiple of the hypercube’s width. See Figure 3. Note that if $A$ is $k$-periodic, then it is also $k^{\prime}$-periodic for every positive integer multiple $k^{\prime}=i\cdot k$ of $k$.
@@
    $\phi(\mathbf{x}_{j})\neq\phi(\mathbf{x}_{j+1})$ (correct answer swaps for each subsequent point),
@@
If a noncollapsing, all-voting, entirely execution bounded CRD stably decides a predicate $\phi$, then $\phi$ is eventually constant.
[…]
submissionA complete proof appears in the appendix.
@@
This proof is similar to that of Theorem 6.9. In that proof, we repeatedly add a “constant amount of additional input $\{X_{2}\}$ or $\{X_{1}\}$, which flips the output”. For more general semilinear, but not eventually constant, predicates, we dig into the structure of the semilinear set to find a sequence of constant-size vectors representing additional inputs that flip the correct answer. Any predicate that is not eventually constant has infinitely many yes inputs and infinitely many no inputs, but in general they could be increasingly far apart: e.g., $\phi(\mathbf{x})=1$ if and only if $2^{n}\leq\|\mathbf{x}\|<2^{n+1}$ for even $n$. For the potential function argument to work, each subsequence input needs to be at most a constant larger than the previous.
@@
    $\phi(\mathbf{x}_{i})\neq\phi(\mathbf{x}_{i+1})$ (i.e., the correct answer swaps for each subsequent input)
@@
    $\mathbf{x}_{i}\leq\mathbf{x}_{i+1}$, i.e., the inputs are increasing (on at least one coordinate(s)), and
@@
    for some constant $c$, $\|\mathbf{x}_{i+1}-\mathbf{x}_{i}\|\leq c$, i.e., adjacent inputs are “close”.
@@
Continuing in this way, we have a sequence of stable configurations $\mathbf{a}_{0},\mathbf{a}_{1},\dots$ where all species in $\mathbf{a}_{i}$ vote yes for odd $i$, and all species in $\mathbf{a}_{i}$ vote no for even $i$. Since $\mathcal{D}$ is noncollapsing, the size of the configurations $\mathbf{a}_{i}$ increases without bound as $i\to\infty$. (Possibly $\|\mathbf{a}_{i+1}\|<\|\mathbf{a}_{i}\|$, i.e., the size is not necessarily monotonically nondecreasing, but for all sufficiently large $j>i$, we have $\|\mathbf{a}_{j}\|>\|\mathbf{a}_{i}\|$.)
@@
However, since $\mathcal{D}$ is noncollapsing, this can only happen for a constant number of $i$ (so $\Phi$ never reaches more than a constant above its initial value $\Phi(\mathbf{x}_{0})$), after which point $\Phi$ strictly decreases after each round of this process.
@@
If a noncollapsing, all-voting, leaderless, execution bounded CRD stably decides a predicate $\phi$, then $\phi$ is eventually constant.
In particular, since the original model of population protocols [angluin2004computation] defined them as leaderless and all-voting—and since population protocols are noncollapsing—we have the following.

### v3

@@ 6 Limitations of execution bounded CRNs
The main positive results of the paper (Theorems 4.1 and 5.5) rely on the assumption that valid initial configurations have a single leader (in particular, they are execution bounded only from configurations with a single leader, but not from arbitrary configurations). Section 4 shows that we may assume the CRD deciding a semilinear set is all-voting. However, for the “constructive” results LABEL:{lem:boolean-closure-crd} and 5.5, which compose the output of a CRD $\mathcal{D}$ with downstream computation, using $\mathcal{D}$ as a “subroutine” to stably compute a more complex set/function, the constructions crucially use the assumption that $\mathcal{D}$ is single-voting (i.e., only the leader of $\mathcal{D}$ votes) to argue the resulting composed CRN is execution bounded. In this section we show these assumptions are necessary, proving that execution bounded CRNs without those constraints are severely limited in their computational abilities.
[…]
@@ 6.1 Linear potential functions
We define a linear potential function of a CRN to be a nonnegative linear function of configurations that each reaction strictly decreases.
@@
A linear potential function $\Phi:\mathbb{R}_{\geq 0}^{\Lambda}\to\mathbb{R}_{\geq 0}$ for a CRN is a nonnegative linear function, such that for each reaction $(\mathbf{r},\mathbf{p})$, $\Phi(\mathbf{p})-\Phi(\mathbf{r})<0.$
Note that for a configuration $\mathbf{x}$, since $\Phi(\mathbf{x})=\sum_{S\in\Lambda}v_{S}\mathbf{x}(S)\geq 0$, it must be nondecreasing in each species, i.e., all coefficients $v_{S}$ must be nonnegative (though some are permitted to be 0). Intuitively, we can think of $\Phi$ as assigning a nonnegative “mass” to each species (the mass of $S$ is $v_{S}$), such that each reaction removes a positive amount of mass from the system. Note also that since $\Phi$ is linear, the above is equivalent to requiring that $\Phi(\mathbf{p}-\mathbf{r})<0$, if we extend $\Phi$ to a linear function $\Phi:\mathbb{R}^{\Lambda}\to\mathbb{R}$ on vectors with negative elements.
[…]
$(\forall i)\sum_{S\in\Lambda}[\mathbf{p}_{i}(S)-\mathbf{r}_{i}(S)]v_{S}<0$
[…]
submission,final $(\forall i)\sum_{S\in\Lambda}[\mathbf{p}_{i}(S)-\mathbf{r}_{i}(S)]v_{S}<0.$ For example, for the reactions $A+A\mathop{\rightarrow}\limits B+C$ and $B+B\mathop{\rightarrow}\limits A$, for each reaction to strictly decrease the potential function $\Phi(\mathbf{x})=v_{A}\mathbf{x}(A)+v_{B}\mathbf{x}(B)+v_{C}\mathbf{x}(C)$, $\Phi$ must satisfy $2v_{A}>v_{B}+v_{C}$ and $2v_{B}>v_{A}$. In this case, $v_{A}=1,v_{B}=1,v_{C}=0$ works.
@@
A system of linear inequalities with rational coefficients has a real solution if and only if it has a rational solution. For any homogeneous system (where all inequalities are comparing to 0), any positive scalar multiple of a solution is also a solution. By clearing denominators, a system has a rational solution if and only if it has an integer solution. Thus, one can equivalently define a linear potential function to be a function $\Phi(\mathbf{x})=\sum_{S\in\Lambda}v_{S}\mathbf{x}(S)$ such that each $v_{S}\in\mathbb{N}$, i.e., we may assume $\Phi:\mathbb{N}^{\Lambda}\to\mathbb{N}$. In particular, since $\Phi$ is decreased by each reaction, it is decreased by at least 1.
@@
full,submission
The following is a variant of Farkas’ Lemma [8], one of several similar “Theorems of the Alternative” stating that exactly one of two different linear systems has a solution. (See [14, Chapter 4] for a list of such theorems.) A proof can be found in [9, Theorem 2.10].
@@
Although we do not need the following fact, it is worthwhile to observe that, if $\mathbf{M}$ is integer-valued (as in our application), then the solution $\mathbf{u}$ or $\mathbf{v}$ (whichever exists) in Section 6.1 has entries that are at most exponential in $\|\mathbf{M}\|$ i.e., at most exponential in the sum of absolute values of entries of $\mathbf{M}$ (see e.g., [15]). So in particular when we consider $\mathbf{M}$ having small $O(1)$ size entries, this means the solution $\mathbf{u}$ or $\mathbf{v}$ has entries that are at most exponential in the number of rows and columns of $\mathbf{M}$. When $\mathbf{M}$ is a stoichiometric matrix, this corresponds to the number of species and reactions, respectively, of the CRN.
@@
[…]
finalThe following theorem due to Czerner, Guttenberg, Helfrich, and Esparza, is crucial to proving limitations on execution bounded CRNs such as Theorems 6.9 and 6.3.
Theorem 6.5 ([6]).
@@
Let $\mathcal{C}=(\Lambda,R)$ be a CRN. The forward direction is easy: assuming $\mathcal{C}$ has potential function $\Phi$, since each reaction decreases $\Phi$ by at least 1 (see Remark 6.2), from any configuration $\mathbf{x}$, $\mathcal{C}$ can execute at most $\Phi(\mathbf{x})$ reactions while keeping $\Phi$ nonnegative. Thus $\mathcal{C}$ is entirely execution bounded.
@@
By Section 6.1, there is an integer vector $\mathbf{v}\geqq\mathbf{0}$ such that $\mathbf{v}\mathbf{M}<\mathbf{0}$. Let $\mathbf{v}\in\mathbb{N}^{\Lambda}$ be the coefficients of a linear function $\Phi:\mathbb{N}^{\Lambda}\to\mathbb{N}$, i.e., $\Phi(\mathbf{x})=\mathbf{v}\cdot\mathbf{x}$. Then the vector $\mathbf{v}\mathbf{M}\in\mathbb{Z}^{R}$ represents the amount $\Phi$ changes by one unit of each reaction, i.e., $\mathbf{v}\mathbf{M}(\alpha)$ is the amount $\Phi$ increases when executing reaction $\alpha$ once. Since $\mathbf{v}\mathbf{M}<\mathbf{0}$, this means that every reaction strictly decreases $\Phi$, i.e., $\Phi$ is a linear potential function for $\mathcal{C}$.
@@
By employing the real-valued version of Section 6.1, the above proof also shows that Theorem 6.5 holds for the continuous model of CRNs [4], in which species amounts are modeled as continuous nonnegative real concentrations. In this case, a continuous CRN would be defined to be execution bounded from configuration $\mathbf{x}$ if each reaction can be executed by at most a finite (real-valued) amount from $\mathbf{x}$.
@@ 6.2 Impossibility of stably deciding majority and parity
No non-collapsing, all-voting, entirely execution bounded CRD can stably decide the majority predicate $[X_{1}\geq X_{2}?]$ or the parity predicate $[X\equiv 1\mod 2?]$.
@@
Continuing in this way, we have a sequence of stable configurations $\mathbf{a}_{1},\mathbf{b}_{1},\mathbf{a}_{2},\mathbf{b}_{2},\dots$ where all species in $\mathbf{a}_{i}$ vote yes and all species in $\mathbf{b}_{i}$ vote no. Since $\mathcal{D}$ is non-collapsing, the size of the configurations $\mathbf{a}_{i}$ and $\mathbf{b}_{i}$ increases without bound as $i\to\infty$. (Possibly $\|\mathbf{a}_{i+1}\|<\|\mathbf{a}_{i}\|$, i.e., the size is not necessarily monotonically increasing, but for all sufficiently large $j>i$, we have $\|\mathbf{a}_{j}\|>\|\mathbf{a}_{i}\|$.) Since all species vote, for some constant $\delta>0$, to get from $\mathbf{a}_{i}+\{X_{2}\}$ to $\mathbf{b}_{i}$, at least $\delta\|\mathbf{a}_{i}\|$ reactions must occur. This is because all species in $\mathbf{a}_{i}$ must be removed since they vote yes, and each reaction removes at most $O(1)$ molecules. (Concretely, let $\delta=1/\max_{(\mathbf{r},\mathbf{p})\in R}\|\mathbf{r}\|-\|\mathbf{p}\|$, i.e., 1 over the most net molecules consumed in any reaction.) Similarly, to get from $\mathbf{b}_{i}+\{X_{1}\}$ to $\mathbf{a}_{i+1}$, at least $\delta\|\mathbf{b}_{i}\|$ reactions must occur.
Since $\mathcal{D}$ is entirely execution bounded, by Theorem 6.5, $\mathcal{D}$ has a linear potential function $\Phi(\mathbf{x})=\mathbf{v}\cdot\mathbf{x}$, where $\mathbf{v}\geq\mathbf{0}$. Adding a single $X_{2}$ to $\mathbf{a}_{i}$ increases $\Phi$ by the constant $\mathbf{v}(X_{2})$. Since $\|\mathbf{a}_{i}\|$ grows without bound, the number of reactions to get from $\mathbf{a}_{i}+\{X_{2}\}$ to $\mathbf{b}_{i}$ increases without bound as $i\to\infty$, and since each reaction strictly decreases $\Phi$ by at least 1, the total change in $\Phi$ that results from adding $X_{2}$ and then going from $\mathbf{a}_{i}+\{X_{2}\}$ to $\mathbf{b}_{i}$ is unbounded in $i$, so unboundedly negative for sufficiently large $i$ (negative once $i$ is large enough that $\delta\|\mathbf{a}_{i}\|\geq\mathbf{v}(X_{2})+2$). Similarly, adding a single $X_{1}$ to $\mathbf{b}_{i}$ and going from $\mathbf{b}_{i}+\{X_{1}\}$ to $\mathbf{a}_{i+1}$, the resulting total change in $\Phi$ is unbounded and (for large enough $i$) negative.
@@
However, since $\mathcal{D}$ is non-collapsing, this can only happen for a constant number of $i$ (so $\Phi$ never reaches more than a constant above its initial value $\Phi(\mathbf{s}+\{X_{1},X_{2}\})$), after which $\Phi$ strictly decreases after each round of this process. At some point in this process, $\mathcal{D}$ will not be able to reach all the way to the next $\mathbf{a}_{i}$ or $\mathbf{b}_{i}$ without $\Phi$ becoming negative, a contradiction.
@@
Theorem 6.9 is false without the non-collapsing hypothesis. The following collapsing, leaderless (but all-voting and entirely execution bounded) CRD stably decides majority: Species $X_{1},x_{1}$ vote yes, while $X_{2},x_{2}$ vote no:
@@
It has bounded executions from every configuration: $\min(\#X_{1},\#X_{2})$ of the first reaction can occur, and the other reactions decrease molecular count, so are limited by the total configuration size. However, it is collapsing since, for any $n$, there exists an input of size $n$ that reaches a stable configuration of size 1. Theorem 6.9 is similarly false without the all-voting hypothesis; for each of the reactions with one product above, add another non-voting product $W$. This converts the CRD to be non-collapsing but not all-voting. Of course, the execution bounded hypothesis is also necessary: the original population protocols paper [1] showed that all-voting, non-collapsing, leaderless population protocols can stably decide all semilinear predicates.
@@
full It has bounded executions from any configuration: exactly #$\#X_{1}+\#X_{0}-1$ reactions can occur since each reduces #$\#X_{1}+\#X_{0}$ by 1. Similar to above, by adding the non-voting product $W$ to each reaction above, this CRD becomes non-collapsing but not all-voting, showing that the all-voting hypothesis is also necessary for stably deciding parity.
@@ 6.3 Impossibility of stably deciding not eventually constant predicates
We now present our main negative result, Section 6.3, which generalizes Theorem 6.9 to show that such CRNs can stably decide only very limited (eventually constant) predicates.
@@
In other words, $A$ is periodic if it is a union of copies of a finite subset $F$ of the $k\times k\times\dots\times k$ hypercube with a corner at the origin, translated in each direction by every nonnegative integer multiple of the hypercube’s width $k$. See Figure 3. Note that if $A$ is $k$-periodic, then it is also $k^{\prime}$-periodic for every positive integer multiple $k^{\prime}=i\cdot k$ of $k$.
@@
    $\phi(\mathbf{x}_{j})\neq\phi(\mathbf{x}_{j+1})$ (correct answer swaps for each subsequent input),
@@
If a non-collapsing, all-voting, entirely execution bounded CRD stably decides a predicate $\phi$, then $\phi$ is eventually constant.
@@
This proof is similar to that of Theorem 6.9. In that proof, we repeatedly add a “constant amount of additional input $\{X_{2}\}$ or $\{X_{1}\}$, which flips the output”. For more general semilinear, but not eventually constant, predicates, we dig into the structure of the semilinear set to find a sequence of constant-size vectors representing additional inputs that flip the correct output. Any predicate that is not eventually constant has infinitely many yes inputs and infinitely many no inputs, but in general they could be increasingly far apart: e.g., $\phi(\mathbf{x})=1$ if and only if $2^{n}\leq\|\mathbf{x}\|<2^{n+1}$ for even $n$. For the potential function argument to work, each subsequent input needs to be at most a constant larger than the previous.
@@
    $\phi(\mathbf{x}_{i})\neq\phi(\mathbf{x}_{i+1})$ (correct answer swaps for each subsequent input),
@@
    $\mathbf{x}_{i}\leq\mathbf{x}_{i+1}$, (inputs are increasing (on at least one coordinate(s)), and
@@
    for some constant $c$, $\|\mathbf{x}_{i+1}-\mathbf{x}_{i}\|\leq c$. (adjacent inputs are “close”)
@@
Continuing in this way, we have a sequence of stable configurations $\mathbf{a}_{0},\mathbf{a}_{1},\dots$ where all species in $\mathbf{a}_{i}$ vote yes for odd $i$, and all species in $\mathbf{a}_{i}$ vote no for even $i$. Since $\mathcal{D}$ is non-collapsing, the size of the configurations $\mathbf{a}_{i}$ increases without bound as $i\to\infty$. (Possibly $\|\mathbf{a}_{i+1}\|<\|\mathbf{a}_{i}\|$, i.e., the size is not necessarily monotonically nondecreasing, but for all sufficiently large $j>i$, we have $\|\mathbf{a}_{j}\|>\|\mathbf{a}_{i}\|$.)
@@
However, since $\mathcal{D}$ is non-collapsing, this can only happen for a constant number of $i$ (so $\Phi$ never reaches more than a constant above its initial value $\Phi(\mathbf{x}_{0})$), after which point $\Phi$ strictly decreases after each round of this process.
@@
If a non-collapsing, all-voting, leaderless, execution bounded CRD stably decides a predicate $\phi$, then $\phi$ is eventually constant.
In particular, since the original model of population protocols [1] defined them as leaderless and all-voting—and since population protocols are non-collapsing—we have the following.